OpenStudy (anonymous):

A rectangular deck measures 15 ft. by 20 ft. What is the length of a diagonal between opposite corners?

OpenStudy (anonymous):

sqrt(15^2 + 20^2) = 25

OpenStudy (anonymous):

use Pythagoras theorem which states  a^2+b^2=c^2

OpenStudy (anonymous):

wow thanks!!

Directrix (directrix):

Think Pythagorean Triples. (15, 20, ?) is a 5 multiple of the triple (3, 4, 5). 3*5 = 15 and 4*5 = 20. So 5*5 = 25 is the missing length (hypotenuse)

OpenStudy (anonymous):

ohhh okay thank you! Very helpful!